OpenStudy (anonymous):

2. The volume of a cube is 512 cm3. What would change if the volume of the cube were measured in cubic meters? A. The volume of the prism would decrease. B. The volume of the prism would increase. C. The number of cubic units would increase. D. The number of cubic units would decrease.

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

because changing the units of measurement does not change the actual volume of the cube

OpenStudy (anonymous):

however when changing the units, the number would be in meters cubed and so it would be a smaller number

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

thank very much

OpenStudy (unklerhaukus):

1 m = 100 cm (1 m)^2 = (100 cm)^2 =100^2 cm^2=10^4 cm^2 (1 m)^3 = (100 cm)^3 =100^3 cm^3=10^6 cm^3

OpenStudy (unklerhaukus):

512 cm^3 = 512 cm^3 x [(m^3)/(10^6 cm^3)] = 512x10^-6 m^3 = 5.12x10^-4 m^3